# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

### v2

@@ Abstract
Privacy-Preserving Federated Learning (PPFL) allows multiple clients to collaboratively train a deep learning model by submitting hidden model updates. Nonetheless, PPFL is vulnerable to data poisoning attacks due to the distributed training nature of clients. Existing solutions have struggled to improve the performance of cross-silo PPFL in poisoned Non-IID data. To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance in poisoned Non-IID data while effectively resisting data poisoning attacks. Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of tampered data distribution on federated learning. Moreover, we utilize two servers to achieve Byzantine-robust aggregation by secure aggregation protocol, which greatly reduces the impact of malicious clients. Theoretical analyses confirm the convergence of PPFPL, and experimental results on publicly available datasets show that PPFPL is effective for resisting data poisoning attacks with Non-IID conditions.
keywords:
@@
Privacy-preserving, cross-silo, federated learning, data poisoning attacks, poisoned Non-IID data.
journal: Nuclear Physics B
1 Introduction
@@
Federated Learning (FL) is a distributed learning paradigm where each client submits its model updates instead of raw training data. In industrial applications, massive data is distributed across independent organizations governed by strict privacy regulations [1]. To break data silos among organizations without compromising privacy, cross-silo FL provides a viable solution for industrial scenarios [2][3]. Specifically, most clients in cross-silo FL are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities. However, client-submitted model updates are vulnerable to privacy inference attacks, which threats the privacy security of cross-silo FL [4][5]. To mitigate the privacy risks, Differential Privacy (DP)-based [6][7] or Homomorphic Encryption (HE)-based [8] [9] PPFL approaches are proposed. In particular, the DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the accuracy of the global model by introducing noise. Conversely, the HE-based approach offers higher privacy security without sacrificing the global model accuracy, thus making it more suitable for cross-silo FL.
@@
While HE-based works [8] [9] demonstrated their effectiveness in privacy preservation within cross-silo FL, they are susceptible to data poisoning attacks due to the distributed training paradigm [10][11]. Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data, and submit model updates derived from the poisoned data, severely impacting the performance of PPFL [12][13]. Moreover, in PPFL, data poisoning attacks are more concealed due to privacy technology obscuring model updates from malicious clients, which raises the difficulty of defense. To audit obscured model updates, recent works [14][15][16][17] utilize HE and Secure Multi-party Computing (SMC) to identify malicious updates in ciphertext. However, these works overlook the heterogeneity of local training data (i.e., Non-IID) issue in FL, which causes distinct inconsistency of benign updates, thereby disrupting the defense against malicious updates.
To distinguish between benign and malicious updates in the Non-IID data, existing solutions [18][19] incorporate clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase. While these solutions are effective against model poisoning attacks, they are inadequate for countering data poisoning attacks. This limitation stems from their reliance on specialized training techniques designed for Non-IID data (e.g., FedProx[20], FedDyn[21], or FedLC[22], etc.), which enhance the performance of global model in Non-IID scenarios. However, in data poisoning attacks, malicious clients deliberately tamper with the features and labels of their training data, thereby forming poisoned Non-IID data. These manipulations distort the optimization direction of local model training and significantly degrade the effectiveness of the specialized training techniques. Consequently, integrating the above defense solutions [18][19] with such training techniques fails to effectively mitigate data poisoning attacks in poisoned Non-IID data. Therefore, a critical challenge remains: how to improve the performance of cross-silo PPFL in poisoned Non-IID data while effectively resisting data poisoning attacks.
@@
Inspired by prototype learning, several works [23][24] designed prototype aggregation to address the Non-IID issue in federated learning, where prototypes are transmitted between the server and clients. Each prototype represents a class feature that is calculated from the mean representations transformed from the samples belonging to the same class. For instance, the works in [23][24] suggest that in recognizing the “cat”, different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”. By exchanging these prototypes, clients can gain more knowledge about the concept of “cat”. Consequently, prototypes, as feature representations independent of data distribution, inspire our work.
In cases of poisoned Non-IID data, tampered data features and data distribution cannot be repaired since the server cannot manipulate the behavior of malicious clients. This prompts us to pose a question: Is it possible to design a PPFL that leverages prototype learning, so that client-submitted model updates are affected only by tampered data features and not by tampered data distribution, while incorporating a secure aggregation protocol to achieve Byzantine-robust results? To answer this question, we propose Privacy-Preserving Federated Prototype Learning framework, named PPFPL, which is suitable for cross-silo scenarios. Our core idea is to utilize prototype learning to address the issue of tampered data distribution, while designing a secure aggregation protocol across two servers to achieve Byzantine-robust aggregation results without compromising privacy. PPFPL employs two “non-colluding” servers who are “honest but curious” to execute a secure aggregation protocol that utilizes cosine similarity as a method to penalize malicious clients. In addition, HE and SMC are adopted as underlying technologies to add the privacy-preserving mechanism of the secure aggregation protocol. The main contributions of the paper are summarized as follows.
  • 1.
    We are the first to introduce prototype learning into PPFL to address data poisoning attacks in across-silo FL, which eliminates the impact of tampered data distribution for client-submitted model updates while significantly improving the FL performance in poisoned Non-IID data.
  • 2.
@@
    We design a secure aggregation protocol across two servers to aggregate client-submitted prototypes with HE and SMC techniques, which ensures that two servers and malicious clients cannot utilize client-submitted prototypes to invade the privacy of benign clients.
  • 3.
@@
The rest of the paper is organized as follows. Section 2 reviews recent PPFL works against poisoning attacks. Section 3 introduces prototype learning and HE. Section 4 formalizes the system model of PPFPL. Our framework is presented in Section 5. Section 6 provide theoretical analysis. Section 7 reviews experimental results. Finally, Section 8 concludes this paper.
2 RELATED WORK
@@
Privacy preservation in FL: Although across-silo FL has naturally certain privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of clients. To resist privacy attacks, DP-based and HE-based approaches are proposed to preserve client-submitted model updates. Specifically, DP-based schemes [6][7][25][26] deploy local differential privacy into model updates to ensure privacy without compromising the utility of model updates. Despite their low computation overheads, these schemes add Gaussian noise or Laplace noise into local model training, which reduces the performance of their global models to some extent. In contrast, HE is a commonly used cryptographic primitive in across-silo PPFL that provides strong privacy preservation without degrading the global model’s accuracy. Specifically, Fang and Qian are one of the first scholars to implement PPFL using HE [8]. They proposed a multi-party machine learning scheme using Paillier [27] technique without compromising participants’ private data. Considering the heavy communication overhead of Paillier, the work in [9] proposed a privacy-preserving federated learning scheme using CKKS (i.e., Cheon-Kim-Kim-Song). This scheme significantly reduces computational overhead associated with ciphertexts. This is because CKKS is more efficient and better suitable to handle large-scale vector and multi-parameter network models than Paillier [28]. However, the aforementioned schemes overlook the threat of data poisoning attacks caused by distributed training.
@@
Resisting data poisoning attacks in FL: Tolpegin et al. demonstrated that data poisoning attacks can significantly reduce the classification accuracy of the global model, even with a small percentage of malicious clients [29]. Additionally, they proposed an aggregated defense strategy that can identify malicious clients in FL to circumvent data poisoning attacks. In addition, Zhang et al. proposed an effective framework capable of tracking the attack time, objective, type, and poisoned location for model updates [30]. Differently, Doku et al. employed SVM to audit client’s local training data for excluding malicious clients, which prevents tampered data from poisoning the global model [31]. However, this violates the privacy of clients to some extent. Furthermore, in practice, the threats of privacy and data poisoning attacks usually coexist in across-silo FL. PPFL approaches can utilize cryptographic primitives to provide privacy preservation for clients, while also concealing data poisoning attacks from malicious clients.
@@
Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15] [14][16][17] combine HE and SMC to identify anomalous model updates in ciphertexts. For instance, the work in [15] proposed a privacy-enhanced FL framework that adopts HE as the underlying technology and provides two servers with a channel to punish malicious clients via gradient extraction of logarithmic function. Similarly, the work in [14] designed a validity checking protocol for ciphertexts under two servers to protect data privacy and adaptively adjust the weight of clients’ gradients to weaken data poisoning attacks. However, these schemes [15] [14][16][17] ignore a fundamental problem, i.e., they neglect the deviations from model updates caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious updates or Non-IID data.
@@
To audit malicious updates in PPFL with Non-IID data, the works in [18][19] are proposed to eliminate the deviations caused by Non-IID data during the aggregation stage. Specifically, ShieldFL designs a Byzantine-tolerant aggregation mechanism to prevent misjudgments on outliers caused by Non-IID data [18]. Furthermore, Chen et al. adopts clustering combined with cosine similarity and median strategies to eliminate deviations among model updates during aggregation auditing [19]. These schemes can only resist model poisoning attacks confronted by federated learning with Non-IID data, but they cannot essentially improve the performance of FL on Non-IID data. They need to be combined with specialized training techniques designed for Non-IID data (e.g., FedProx[20], FedDyn[21], or FedLC[22], etc.) to radically improve performance of FL. These specialized techniques adds an auxiliary term into local loss function to constrain model updates of clients, which helps to increase consistency of model updates during local model training. However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, creating poisoned Non-IID data, which causes specialized training techniques to constrain model updates based on these compromised inputs. Therefore, the above works [18][19] cannot be combined with the specialized techniques to resist data poisoning attacks while improving the performance of the global model under poisoned Non-IID data.
@@
Federated prototype learning: Recently, prototype learning is gradually being applied in federated learning to solve the Non-IID issue. Specifically, the works in [23][24][32] are one of the first to propose federated prototype learning using the concept of prototype learning. Different from the specialized training techniques (e.g., FedProx, FedDyn, or FedLC), its core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes. In other words, each class holds its corresponding prototype that is independent of other classes. As a result, client-submitted prototype is affected by the samples and is independent of data distribution among clients. This inspires our work: federated prototype learning allows distributed training to remain independent of tampered data distributions caused by data poisoning attacks.
@@
Although prototype learning has already been studied to improve the performance of federated learning in Non-IID data, it has not been explored in privacy-preserving federated learning against data poisoning attacks. Our work is the first to integrate prototype learning into PPFL to tackle the performance degradation in across-silo FL caused by tampered Non-IID data, while preserving the privacy of client-submitted prototypes.
3 PRELIMINARIES
@@
3.1 Prototypes Meet Federated Learning
In the classification task of prototype learning, the prototype is defined as a feature vector representing a specific class [33]. Due to this property, the prototypes of the same class among clients are similar in FL task. Consequently, many FL schemes [23][32] enhance the handling of Non-IID data by prototype learning, which allows clients to align their prototypes with other clients during local model training.
@@ 4.1 PPFPL framework
Fig. 1: The PPFPL framework. ① KGC generates $Pkv/Skv$ for Verifier and $Pkx/Skx$ for Clients. ② After local training is completed, clients submit encrypted local prototypes to the Aggregator. ③ Verifier and Aggregator perform secure aggregation protocol to get encrypted global prototypes. ④ Aggregator distributes encrypted global prototypes to Clients.
  • 1.
@@
4.2 Potential Threats for PPFPL
@@
4.3 Design Goals of PPFPL
@@ 5.1 Overview of PPFPL
  • 1.
    Step I. Local Computation: Each client $m$ trains its local model with its local dataset $\mathcal{D}_{m}$. Then, the client normalizes and encrypts its local prototypes and submits them to Aggregator.
  • 2.
@@ 13       Aggregator distributes global prototypes to each client;
5.2 Local Computation
The local computation step includes two essential stages: local model training and prototype handling. The detail is outlined in Algorithm 2.
@@ 5.2 Local Computation
5.2.1 Local Model Training
During local model training, each client aims to minimize classification loss while aligning its local prototype close to the global prototype. To achieve this, an auxiliary term is incorporated into the local loss function. Formally, the local loss function of client $m$ is defined as:
@@ 5.2.1 Local Model Training
where $\textit{sim}(\cdot,\cdot)$ denotes cosine similarity between two vectors. The cosine similarity ranges from -1 to 1, where the value closer to “1” indicates similar vector directions, and conversely, the value closer to “-1” means opposing directions. The objective of each client aims to minimize classification loss while keeping its local prototype close to the global prototype in the direction. Next, we present the process of local model training.
Specifically, in the $t$-th communication round, each client $m$ uses its local dataset $\mathcal{D}_{m}$ to iteratively train its local model. This iterative process is executed in Algorithm 3. The superscript $(e)$ indicates the iteration state of variables, where $e\in\{1,\cdots,E\}$. Firstly, each client $m$ uses the local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the $(t-1)$-th round as the starting point in the current $t$-th round. Subsequently, each client iteratively performs the following stages.
  • 1.

### v4

@@ Abstract
Privacy-Preserving Federated Learning (PPFL) allows multiple clients to collaboratively train by submitting hidden model updates. Nonetheless, PPFL is vulnerable to data poisoning attacks due to the distributed training nature of clients in cross-silo scenarios. Existing solutions have struggled to improve the performance of PPFL in poisoned Non-IID data. To address the issues, this paper proposes a privacy-preserving federated prototype learning framework, named PPFPL, which enhances the cross-silo FL performance in poisoned Non-IID data while protecting the privacy of clients. Specifically, we adopt prototypes as client-submitted model updates to eliminate the impact of poisoned data distribution. In addition, we design a secure aggregation protocol by utilizing homomorphic encryption to achieve Byzantine-robust aggregation on two servers, which greatly reduces the impact of malicious clients. Theoretical analyses confirm the convergence and privacy of PPFPL, and experimental results on publicly datasets show that PPFPL is effective for resisting data poisoning attacks with Non-IID data.
Index Terms:
@@
Privacy-Preserving, Federated Learning, Cross-Silo, Data Poisoning Attacks, Poisoned Non-IID Data.
I Introduction
@@
Federated Learning (FL) is a distributed learning paradigm where each client shares its model updates instead of raw training data. In industrial applications, massive data is distributed across independent organizations governed by strict privacy regulations [1]. To break data silos among organizations without compromising privacy, cross-silo FL provides a viable solution for industrial scenarios [2][3]. Specifically, in cross-silo FL, the clients are usually large organizations, resulting in a relatively small number of clients with significant computational capabilities. However, client-submitted model updates are vulnerable to privacy attacks, which threats the privacy security of cross-silo FL [4][5]. To mitigate the privacy risks, Differential Privacy (DP)-based [6][7] and Homomorphic Encryption (HE)-based [8] [9] PPFL approaches are proposed. In particular, DP-based approach is commonly applied to cross-device FL due to its low computation overhead, but degrades the FL performance by introducing noise. Conversely, HE-based approach offers higher privacy security without sacrificing the global model accuracy, thus making it more suitable for cross-silo FL.
@@
While HE-based approaches demonstrated their effectiveness in terms of privacy preservation, they are susceptible to data poisoning attacks [10][11]. Specifically, malicious clients launch data poisoning attacks by tampering with their raw training data, and submit model updates derived from the poisoned data, thereby degrading the performance of PPFL [12][13]. Moreover, Moreover, privacy-preserving techniques obscure the model updates from malicious clients, making data poisoning attacks more concealed and more difficult to defend. To audit obscured model updates, existing methods [14, 15, 16, 17] utilize HE and Secure Multi-party Computing (SMC) to identify malicious model updates in ciphertext. However, these works overlook the data heterogeneity in FL, i.e., the data among clients is typically non-independent and identically distributed (Non-IID). The Non-IID data leads to the inconsistency among benign model updates, making it difficult for these defense methods [14, 15, 16, 17] to distinguish whether the deviations are caused by data heterogeneity or by malicious model updates. To distinguish between benign and malicious updates in the Non-IID data, existing methods [18, 19, 20] employ clustering or adaptive aggregation weighting operation to mitigate the inconsistency of model updates during the auditing phase, thereby filtering out updates caused by model poisoning attacks. However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, thereby forming poisoned Non-IID data. These manipulations distort the optimization direction of local training, causing these methods [18, 19, 20] to optimize toward the tampered Non-IID distribution, thereby degrading the performance of FL. Therefore, a critical challenge remains: how to enhance the performance of PPFL in poisoned Non-IID data while effectively resisting data poisoning attacks.
@@
Inspired by prototype learning [21], several works [22, 23, 24, 25] proposed prototype aggregation to address the Non-IID issue in FL, where prototypes are exchanged between the server and clients. Each prototype represents a class-level feature, computed as the mean of the feature representations of samples belonging to the same class. For instance, the work in [23] suggest that in recognizing the “cat”, different clients have their own unique “imaginary picture” or “prototype” to represent the concept of “cat”. By exchanging these prototypes, clients gain more knowledge about the concept of “cat”. Consequently, prototypes, as feature representations independent of data distribution, inspire our work.
[…]
Our main contributions are summarized as follows.
  • We are the first to introduce prototype learning into PPFL to defend against data poisoning attacks. By transferring prototypes between clients and servers, our framework mitigates the impact of tampered data distribution and enhances the FL performance in poisoned Non-IID data.
@@
  • We utilize HE and SMC techniques to design a secure aggregation protocol across two servers to filter prototypes submitted by malicious clients, ensuring Byzantine-robust results while protecting the privacy of benign clients.
@@
The rest of the paper is organized as follows. Section II reviews PPFL works against poisoning attacks. Section III introduces prototype learning and HE. Section IV formalizes the system model. Our PPFPL is detailed in Section V. Section VI provide theoretical analysis. Section VII reviews experimental results. Finally, Section VIII concludes this paper.
II RELATED WORK
@@
Privacy preservation in FL: Although FL has naturally certain privacy protection, it remains vulnerable to privacy attacks, which causes the privacy threat of benign clients. To resist such attacks, DP-based and HE-based approaches are proposed to preserve client-submitted model updates. Specifically, DP-based schemes [26, 27, 28, 29] deploy local differential privacy into model updates, ensuring privacy without compromising the utility of model updates. Despite their low computation overheads, these schemes add Gaussian noise or Laplace noise into local model training, which degrades the FL performance to some extent. In contrast, HE is a commonly used cryptographic primitive in across-silo PPFL that provides strong privacy preservation without sacrificing the global model’s accuracy. Specifically, Fang and Qian are one of the first scholars to implement PPFL using HE [30]. They proposed a multi-party machine learning scheme using Paillier [31] technique without compromising clients’ privacy. Considering the heavy communication overhead of Paillier, the work in [32] proposed a privacy-preserving FL using CKKS (i.e., Cheon-Kim-Kim-Song) that reduces computational overhead associated with ciphertexts. This is because CKKS is more efficient and better suitable to handle large-scale vector and multi-parameter network models than Paillier [33]. However, the aforementioned schemes overlook the threat of data poisoning attacks caused by distributed training.
@@
Resisting data poisoning attacks in FL: Tolpegin et al. demonstrated that data poisoning attacks can significantly reduce the classification accuracy of the global model, even with a small percentage of malicious clients [34]. Additionally, they proposed an aggregated defense strategy that identifies malicious clients in FL to circumvent data poisoning attacks. Similarly, the works in [35, 36, 37] proposed detection methods to filter malicious model updates. Differently, Doku et al. employed SVM to audit client’s local training data for excluding malicious clients [38]. However, this work violates the privacy of clients to some extent. Furthermore, the threats of privacy and data poisoning attacks usually coexist in practice across-silo FL. PPFL approaches can utilize cryptographic primitives to provide privacy preservation for clients, while concealing data poisoning attacks from malicious clients.
@@
Resisting data poisoning attacks in PPFL: Considering the threats of both privacy and data poisoning attacks, the works such as [15, 14, 16, 17] combine HE and SMC to identify anomalous model updates in ciphertexts. For instance, the work in [15] proposed a privacy-enhanced FL framework that adopts HE as the underlying technology and provides two servers with a channel to punish malicious clients via gradient extraction of logarithmic function. Similarly, the work in [14] designed a validity checking protocol for ciphertexts under two servers to protect data privacy and adaptively adjust the weight of clients’ gradients to weaken data poisoning attacks. However, these schemes [15, 14, 16, 17] ignore a fundamental problem, i.e., they neglect the deviations from model updates caused by Non-IID data, making it difficult to distinguish whether the deviations come from malicious model updates or Non-IID data.
@@
To audit malicious updates in PPFL with Non-IID data, the works in [18][19] are proposed to eliminate the deviations caused by Non-IID data during the aggregation stage. Specifically, ShieldFL designs a Byzantine-tolerant aggregation mechanism to prevent misjudgments on outliers caused by Non-IID data [18]. Furthermore, Chen et al. adopts clustering combined with cosine similarity and median strategies to eliminate deviations among model updates during aggregation auditing [19]. These schemes can only resist model poisoning attacks confronted by federated learning with Non-IID data, but they cannot essentially improve the performance of FL on Non-IID data. They need to be combined with specialized training techniques designed for Non-IID data (e.g., FedProx[39], FedDyn[40], or FedLC[41], etc.) to radically improve performance of FL. These specialized techniques adds an auxiliary term into local optimization function to constrain model updates of clients, which helps to increase consistency of model updates during local model training. However, in data poisoning attacks, malicious clients tamper with the features and labels of their training data, creating poisoned Non-IID data, which causes specialized training techniques to constrain model updates based on these compromised inputs. Therefore, the above works [18][19] cannot be combined with the specialized techniques to resist data poisoning attacks while enhancing the FL performance under poisoned Non-IID data.
@@
Federated prototype learning: Recently, prototype learning is gradually being applied in federated learning to solve the Non-IID issue. Specifically, the works in [22, 23, 24, 25] are one of the first to propose federated prototype learning using the concept of prototype learning. Different from the specialized training techniques (e.g., FedProx, FedDyn, or FedLC), the core idea enables clients to pull the same-class samples towards the global prototypes of that class and away from the global prototypes of other classes. In other words, each class holds its corresponding prototype that is independent of other classes. As a result, client-submitted prototype is affected by the samples but is independent of data distribution among clients. This inspires our work: federated prototype learning allows distributed training to remain independent of tampered data distributions caused by data poisoning attacks.
@@
Although prototype learning has already been studied to improve the performance of federated learning in Non-IID data, it has not been explored in privacy-preserving federated learning against data poisoning attacks. Our work is the first to integrate prototype learning into PPFL to tackle the performance degradation caused by tampered Non-IID data in across-silo FL, while preserving the privacy of client-submitted prototypes.
III PRELIMINARIES
@@
III-A Prototypes Meet Federated Learning
In the classification task of prototype learning, the prototype is defined as a feature vector representing a specific class [42]. Due to this property, the prototypes of the same class among clients are similar in FL task. Consequently, many FL schemes [22, 23, 24, 25] enhance the handling of Non-IID data by prototype learning, which allows clients to align their prototypes with other clients during local model training.
@@ IV-A PPFPL framework
Figure 1: The PPFPL framework. ① KGC generates $Pkv/Skv$ for Verifier and $Pkx/Skx$ for Clients. ② After local training is completed, clients submit encrypted local prototypes to the Aggregator. ③ Verifier and Aggregator perform secure aggregation protocol to get encrypted global prototypes. ④ Aggregator distributes encrypted global prototypes to Clients.
@@
IV-B Potential Threats for PPFPL
@@
IV-C Design Goals of PPFPL
@@ V-A Overview of PPFPL
  • Step I. Local Computation: Each client trains its local model with its local dataset. Then, the client normalizes and encrypts its local prototypes and submits them to Aggregator.
@@ 13       Aggregator distributes global prototypes to each client;
V-B Local Computation
The local computation step includes two essential stages: local model training and prototype handling stages. The detail is outlined in Algorithm 2.
@@ V-B Local Computation
V-B1 Local Model Training
During local model training, each client aims to minimize classification loss while aligning its local prototype close to the global prototype. To achieve this, we design an auxiliary term by leveraging prototype learning in the local optimization function. Formally, the local optimization function of client $m$ is defined as:
@@ V-B1 Local Model Training
where $\textit{sim}(\cdot,\cdot)$ denotes cosine similarity between two vectors. The cosine similarity ranges from -1 to 1, where the value closer to “1” indicates similar vector directions, and conversely, the value closer to “-1” means opposing directions. The optimization function ensures that each client reduces the classification loss while aligning its local prototype with the global prototype in direction. Next, we present the process of local model training.
Specifically, in the $t$-th communication round, each client $m$ uses its local dataset $\mathcal{D}_{m}$ to iteratively train its local model. This iterative process is executed in Algorithm 3. The superscript $(e)$ indicates the iteration state of variables, where $e\in\{1,\cdots,E\}$. Each client $m$ uses its local model parameters $\textit{{w}}_{m,t-1}^{(E)}$ from the $(t-1)$-th round as the starting point in the current $t$-th round. Subsequently, each client iteratively performs the following stages.